ROTATIONAL SPECTROSCOPY AND MOLECULAR PROPERTIES OF TITANIUM DIOXIDE, TiO_{2}

\underline{HOLGER S. P. MÜLLER}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln; and Max-Planck-Institut für Radioastronomie, 53121 Bonn, Germany; SANDRA BRÜNKEN, MICHAEL C. McCARTHY, PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138, USA.

Titanium is one of the more abundant transition metals in space, and its dioxide is thought to be an important constituent of grains generated in the vicinity of oxygen-rich stars. Therefore, we have investigated the rotational spectrum of TiO_{2} by laser-ablation molecular-beam Fourier transform microwave (LAMB-FTMW) spectroscopy between 7 and 42 GHz. Five isotopic species containing ^{46−50}Ti were studied in natural isotopic composition and two, ^{48}Ti^{16}O^{18}O and ^{48}Ti^{18}O_{2} were obtained through the use of ^{18}O_{2}. Despite the comparatively large rotational constants and the absence of half of the rotational levels for all but the mixed ^{16}O^{18}O isotopic species because of spin-statistics, up to 13 rotational transitions with energies up to almost 40 cm^{−1} could be recorded, permitting rotational and quartic centrifugal distortion parameters to be determined along with hyperfine structure parameters for the species containing ^{47}Ti or ^{49}Ti. The derived molecular structure and the harmonic force field as well as the hyperfine structure parameters will be compared with those of related molecules. Preliminary results of a search toward several oxygen-rich late-type stars carried out with the IRAM 30 m telescope will also be presented.